I havne't taken an algebra class in over twenty years and need help solving two problem:
2. Bob Brown bought two plots of land for a total of $110,000. On the first plot, he made a profit of 16%. On the second, he lost 4%. His total profit was $9600. How much did he pay for each piece of land?
3. A major car rental firm charges $57 a day with unlimited mileage. A discount firm offers a similar car for $24 a day plus 22 cents per mile. How far must you drive in a day in order for the cost to be the same at both firms?

Answer by checkley79(3341)


.16X-.04(110000-X)=9600
.16X-4400+.04X=9600
.20X=9600+4400
.20X=14000
X=14000/.20
X=$70000 VALUE OF THE LOT @ 16%
110000-70000=40000 VALUE OF LOT @ 4%
PROOF:
.16*70000-.04*40000=9600
11200-1600=9600
9600=9600
###################################
57=24+.22X
.22X=57-24
.22X=33
X=33/.22
X=150 MILES.
PROOF:
57=24+.22*150
57=24+33
57=57